# Well-posedness of stochastic mSQG equations with Kraichnan noise and $L^{p}$ data

### v1

@@ 4 Weak existence
We claim that $\tilde{\theta}$ satisfies the bounds (1.8) and (1.10). Since each $\tilde{u}^{\delta_{n}}$ has the same law as $u^{\delta_{n}}$, there exists a version of $\tilde{u}^{\delta_{n}}$ (still called $\tilde{u}^{\delta_{n}}$) which has Hölder continuous paths with values in $H^{-5+\beta}$ and satisfies the bounds (4.1) and (4.2). Observe that the norms in $C_{t}^{\gamma}(H^{-5+\beta})$ and $L_{t}^{\infty}(\dot{H}^{-1+\frac{\beta}{2}})$ are lower semicontinuous functions in $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)$ (see e.g. [10, Lemma B.2]), the limit $\tilde{u}$ (a version) has trajectories in $L_{t}^{\infty}(\dot{H}^{-1+\frac{\beta}{2}})\cap C_{t}^{\gamma}(H^{-5+\beta})$. The uniform $L^{p}$ bound of trajectories of $\tilde{\theta}$ in (1.8) is verified as follows. For every $t\in(0,T),h>0$ sufficiently small and $\varphi\in C_{c}^{\infty}$ with $\|\varphi\|_{L^{p^{\prime}}}\leq 1$, we have $\tilde{\mathbb{P}}$-a.s.,
@@
Letting $n\rightarrow\infty$, from Lemma 4.3 below, we have $\frac{1}{2h}\int_{t-h}^{t+h}\langle\tilde{\theta}_{s},\varphi\rangle\mathrm{d}s\leq\|\theta_{0}\|_{L^{p}}$. Next let $h\rightarrow 0$ and the time continuity implies that $\langle\tilde{\theta}_{t},\varphi\rangle\leq\|\theta_{0}\|_{L^{p}}$, $\tilde{\mathbb{P}}$-a.s. for all $t\in[0,T]$. Then taking supreme at left-hand side for $\varphi\in C_{c}^{\infty}$ with $\|\varphi\|_{L^{p^{\prime}}}\leq 1$, we have $\sup_{t\in[0,T]}\|\tilde{\theta}_{t}\|_{L^{p}}\leq\|\theta_{0}\|_{L^{p}}$. Hence we have the bound (1.8). The proof of the bound (1.10) is very similar, see also [10, Section 6].
@@
$\begin{split}\langle\theta_{t}^{\delta_{n}},\varphi\rangle&=\langle\theta_{0}^{\delta_{n}},\varphi\rangle+\int_{0}^{t}\langle(K_{\beta}^{\delta_{n}}*\theta_{r}^{\delta_{n}})\theta_{r}^{\delta_{n}},\nabla\varphi\rangle\,\mathrm{d}r+\sum_{k}\int_{0}^{t}\langle\sigma_{k}^{\delta_{n}}\theta_{r}^{\delta_{n}},\nabla\varphi\rangle\,\mathrm{d}B_{r}^{\delta_{n},k}\\ &\quad+\frac{c_{\delta_{n}}}{2}\int_{0}^{t}\langle\theta_{r}^{\delta_{n}},\Delta\varphi\rangle\,\mathrm{d}r\\ &:=A_{1}+A_{2}+A_{3}+A_{4}.\end{split}$ (4.3)
@@
Let $\psi=\nabla^{\perp}(-\Delta)^{-1+\frac{\beta}{2}}\varphi=G_{2-\beta}*(\nabla^{\perp}\varphi)$, then we have
@@
Concerning the convergence of the stochastic term $A_{3}$, we use a classical result (see [14, Lemma 2.1]): if
@@
To cope with the nonlinear term $A_{2}$, recalling that $u^{\delta_{n}}=K_{\beta}*\theta^{\delta_{n}}$, we have
$\begin{split}&\int_{0}^{t}\big{(}\big{\langle}(K_{\beta}^{\delta_{n}}*\theta_{r}^{\delta_{n}})\theta_{r}^{\delta_{n}},\nabla\varphi\big{\rangle}-\big{\langle}u_{r}\theta_{r},\nabla\varphi\big{\rangle}\big{)}\,\mathrm{d}r\\ &=\int_{0}^{t}\big{\langle}((K_{\beta}^{\delta_{n}}-K_{\beta})*\theta_{r}^{\delta_{n}})\theta_{r}^{\delta_{n}},\nabla\varphi\big{\rangle}\,\mathrm{d}r+\int_{0}^{t}\langle(u_{r}^{\delta_{n}}-u_{r})\theta_{r}^{\delta_{n}},\nabla\varphi\rangle\,\mathrm{d}r\\ &\quad+\int_{0}^{t}\langle(\theta_{r}^{\delta_{n}}-\theta_{r})u_{r},\nabla\varphi\rangle\,\mathrm{d}r\\ &=:A_{21}+A_{22}+A_{23}.\end{split}$
@@ 5 Pathwise uniqueness
$\mathrm{d}\theta+\big{[}(K_{\beta}*\theta^{1})\cdot\nabla\theta+(K_{\beta}*\theta)\cdot\nabla\theta^{2}\big{]}\,\mathrm{d}t+\sum_{k}\sigma_{k}\cdot\nabla\theta\,\mathrm{d}B^{k}=\frac{\pi}{4\alpha}\Delta\theta\,\mathrm{d}t.$
@@
$\begin{split}\mathrm{d}\langle\theta,G_{\beta}^{\delta}*\theta\rangle&=2\langle\nabla G_{\beta}^{\delta}*\theta,(K_{\beta}*\theta^{1})\theta\rangle\,\mathrm{d}t+2\langle\nabla G_{\beta}^{\delta}*\theta,(K_{\beta}*\theta)\theta^{2}\rangle\,\mathrm{d}t\\ &\quad+2\sum_{k}\langle\nabla G_{\beta}^{\delta}*\theta,\sigma_{k}\theta\rangle\,\mathrm{d}B^{k}\\ &\quad+\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}(Q(0)-Q(x-y))D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta(x)\theta(y)\,\mathrm{d}x\mathrm{d}y\mathrm{d}t\\ &=:(2I_{1}+2I_{2})\,\mathrm{d}t+2\sum_{k}M_{k}\,\mathrm{d}B^{k}+J\,\mathrm{d}t.\end{split}$ (5.1)
@@
$\begin{split}|I_{1}|&=|\,\langle\nabla G_{\beta}^{\delta}*\theta,(K_{\beta}*\theta^{1})\theta\rangle\,|\\ &\leq\|\theta\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|(K_{\beta}*\theta^{1})\nabla G_{\beta}^{\delta}*\theta\|_{\dot{H}^{\frac{\beta}{2}-1+\alpha+\epsilon}}\\ &\lesssim\|\theta\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|K_{\beta}*\theta^{1}\|_{\dot{H}^{2(\alpha+\epsilon)}}\|{\nabla G_{\beta}^{\delta}*\theta}\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}\\ &\lesssim\|\theta\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}^{2}\|\theta^{1}\|_{\dot{H}^{2(\alpha+\epsilon)-\beta+1}}.\end{split}$
@@
$\langle\nabla G_{\beta}*\theta,(K_{\beta}*\theta)\theta^{2}\rangle=\langle\nabla(G_{\beta}*\theta),\nabla^{\perp}(G_{\beta}*\theta)\,\theta^{2}\rangle=0,$
@@
$\begin{split}|I_{2}|&=|\langle\nabla(G_{\beta}^{\delta}-G_{\beta})*\theta,(K_{\beta}*\theta)\theta^{2}\rangle|\\ &\leq\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|(K_{\beta}*\theta)\nabla(G_{\beta}^{\delta}-G_{\beta})*\theta\|_{\dot{H}^{\frac{\beta}{2}-1+\alpha+\epsilon}}\\ &\lesssim\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|K_{\beta}*\theta\|_{\dot{H}^{2(\alpha+\epsilon)}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})*\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}\\ &\leq\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|\theta\|_{\dot{H}^{2(\alpha+\epsilon)-\beta+1}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})*\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}.\end{split}$
@@
$\begin{split}|I_{2}|&\lesssim\big{(}\|\theta^{1}\|_{L^{\tilde{p}}}+\|\theta^{2}\|_{L^{\tilde{p}}}\big{)}\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})*\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}\\ &\lesssim(C_{1}+C_{2})\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})*\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}.\end{split}$
@@
$\int_{0}^{T}\mathbb{E}[|I_{2}|]\,\mathrm{d}r\lesssim\bigg{(}\int_{0}^{T}\mathbb{E}\Big{[}\|\theta_{r}^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}^{2}\Big{]}\mathrm{d}r\bigg{)}^{1/2}\bigg{(}\int_{0}^{T}\mathbb{E}\Big{[}\|\nabla(G_{\beta}^{\delta}-G_{\beta})*\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}^{2}\Big{]}\mathrm{d}r\bigg{)}^{1/2}.$
@@
For $(t,x)\in(0,\infty)\times\mathbb{R}^{d}$, define $q(t,x)=t(t^{2/\beta}+|x|^{2})^{-\frac{d+\beta}{2}}$. The following bound of the kernel $p(t,x)$ is well known, see for example [5; 31].
@@
From a probabilistic point of view, the fractional Laplacian $-(-\Delta)^{\frac{\beta}{2}}$ is the infinitesimal generator of the $\beta$-stable process, which is a pure jump Markov process. That is to say, $p(t,x)$ is the transition probability of a $\beta$-stable process $Z=(Z_{t})_{t\geq 0}$. Then by the Bochner subordination formula (see e.g. [4, Proposition 8.6]), we have
@@
where $h(s,x)=(4\pi s)^{-\frac{d}{2}}e^{-\frac{|x|^{2}}{4s}}$ is the heat kernel in $\mathbb{R}^{d}$, $\rho_{t}$ is the probability density function of $S_{t}$. Accroding to [5, Lemma 5], there exists a constant $c$ such that $\rho_{1}(s)\leq c\,s^{-1-\frac{\beta}{2}}$ for all $s>0$, so by scaling property $\rho_{t}(s)=t^{-\frac{2}{\beta}}\rho_{1}(t^{-\frac{2}{\beta}}s)$, we have
@@ Appendix B Estimating the remainders
The proofs are almost identical to [10, Lemma A.1], so we omit the details here.
@@
Since $21<p<\frac{2}{3-\beta+\gamma}<$ 2, we can choose $0<\lambda<1-\gamma$ such that $22+(\gamma+\lambda-1)p<$ 2 and $p(4-\beta-\lambda)<2$, making the right-hand side above integrable. Therefore, $g$ has finite $W^{1+\gamma,p}$ norm for $10<\gamma<\beta-$ 1 and $21<p<\frac{2}{3-\beta+\gamma}<$ 2. Then we have (see [40, Subsection 2.5])

### v2

@@ 4 Weak existence
We claim that $\tilde{\theta}$ satisfies the bounds (1.8) and (1.10). Since each $\tilde{u}^{\delta_{n}}$ has the same law as $u^{\delta_{n}}$, there exists a version of $\tilde{u}^{\delta_{n}}$ (still called $\tilde{u}^{\delta_{n}}$) which has Hölder continuous paths with values in $H^{-5+\beta}$ and satisfies the bounds (4.1) and (4.2). Observe that the norms in $C_{t}^{\gamma}(H^{-5+\beta})$ and $L_{t}^{\infty}(\dot{H}^{-1+\frac{\beta}{2}})$ are lower semicontinuous functions in $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)$ (see e.g. [11, Lemma B.2]), the limit $\tilde{u}$ (a version) has trajectories in $L_{t}^{\infty}(\dot{H}^{-1+\frac{\beta}{2}})\cap C_{t}^{\gamma}(H^{-5+\beta})$. The uniform $L^{p}$ bound of trajectories of $\tilde{\theta}$ in (1.8) is verified as follows. For every $t\in(0,T),h>0$ sufficiently small and $\varphi\in C_{c}^{\infty}$ with $\|\varphi\|_{L^{p^{\prime}}}\leq 1$, we have $\tilde{\mathbb{P}}$-a.s.,
@@
Letting $n\rightarrow\infty$, from Lemma 4.3 below, we have $\frac{1}{2h}\int_{t-h}^{t+h}\langle\tilde{\theta}_{s},\varphi\rangle\mathrm{d}s\leq\|\theta_{0}\|_{L^{p}}$. Next let $h\rightarrow 0$ and the time continuity implies that $\langle\tilde{\theta}_{t},\varphi\rangle\leq\|\theta_{0}\|_{L^{p}}$, $\tilde{\mathbb{P}}$-a.s. for all $t\in[0,T]$. Then taking supreme at left-hand side for $\varphi\in C_{c}^{\infty}$ with $\|\varphi\|_{L^{p^{\prime}}}\leq 1$, we have $\sup_{t\in[0,T]}\|\tilde{\theta}_{t}\|_{L^{p}}\leq\|\theta_{0}\|_{L^{p}}$. Hence we have the bound (1.8). The proof of the bound (1.10) is very similar, see also [11, Section 6].
@@
$\begin{split}\langle\theta_{t}^{\delta_{n}},\varphi\rangle&=\langle\theta_{0}^{\delta_{n}},\varphi\rangle+\int_{0}^{t}\langle(K_{\beta}^{\delta_{n}}\ast\theta_{r}^{\delta_{n}})\theta_{r}^{\delta_{n}},\nabla\varphi\rangle\,\mathrm{d}r+\sum_{k}\int_{0}^{t}\langle\sigma_{k}^{\delta_{n}}\theta_{r}^{\delta_{n}},\nabla\varphi\rangle\,\mathrm{d}B_{r}^{\delta_{n},k}\\ &\quad+\frac{c_{\delta_{n}}}{2}\int_{0}^{t}\langle\theta_{r}^{\delta_{n}},\Delta\varphi\rangle\,\mathrm{d}r\\ &:=A_{1}+A_{2}+A_{3}+A_{4}.\end{split}$ (4.3)
@@
Let $\psi=\nabla^{\perp}(-\Delta)^{-1+\frac{\beta}{2}}\varphi=G_{2-\beta}\ast(\nabla^{\perp}\varphi)$, then we have
@@
Concerning the convergence of the stochastic term $A_{3}$, we use a classical result (see [15, Lemma 2.1]): if
@@
To cope with the nonlinear term $A_{2}$, recalling that $u^{\delta_{n}}=K_{\beta}\ast\theta^{\delta_{n}}$, we have
$\begin{split}&\int_{0}^{t}\big{(}\big{\langle}(K_{\beta}^{\delta_{n}}\ast\theta_{r}^{\delta_{n}})\theta_{r}^{\delta_{n}},\nabla\varphi\big{\rangle}-\big{\langle}u_{r}\theta_{r},\nabla\varphi\big{\rangle}\big{)}\,\mathrm{d}r\\ &=\int_{0}^{t}\big{\langle}((K_{\beta}^{\delta_{n}}-K_{\beta})\ast\theta_{r}^{\delta_{n}})\theta_{r}^{\delta_{n}},\nabla\varphi\big{\rangle}\,\mathrm{d}r+\int_{0}^{t}\langle(u_{r}^{\delta_{n}}-u_{r})\theta_{r}^{\delta_{n}},\nabla\varphi\rangle\,\mathrm{d}r\\ &\quad+\int_{0}^{t}\langle(\theta_{r}^{\delta_{n}}-\theta_{r})u_{r},\nabla\varphi\rangle\,\mathrm{d}r\\ &=:A_{21}+A_{22}+A_{23}.\end{split}$
@@ 5 Pathwise uniqueness
$\mathrm{d}\theta+\big{[}(K_{\beta}*\theta^{1})\cdot\nabla\theta+(K_{\beta}\ast\theta)\cdot\nabla\theta^{2}\big{]}\,\mathrm{d}t+\sum_{k}\sigma_{k}\cdot\nabla\theta\,\mathrm{d}B^{k}=\frac{\pi}{4\alpha}\Delta\theta\,\mathrm{d}t.$
@@
$\begin{split}\mathrm{d}\langle\theta,G_{\beta}^{\delta}\ast\theta\rangle&=2\langle\nabla G_{\beta}^{\delta}\ast\theta,(K_{\beta}\ast\theta^{1})\theta\rangle\,\mathrm{d}t+2\langle\nabla G_{\beta}^{\delta}\ast\theta,(K_{\beta}\ast\theta)\theta^{2}\rangle\,\mathrm{d}t\\ &\quad+2\sum_{k}\langle\nabla G_{\beta}^{\delta}\ast\theta,\sigma_{k}\theta\rangle\,\mathrm{d}B^{k}\\ &\quad+\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}(Q(0)-Q(x-y))D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta(x)\theta(y)\,\mathrm{d}x\mathrm{d}y\mathrm{d}t\\ &=:(2I_{1}+2I_{2})\,\mathrm{d}t+2\sum_{k}M_{k}\,\mathrm{d}B^{k}+J\,\mathrm{d}t.\end{split}$ (5.1)
@@
$\begin{split}|I_{1}|&=|\,\langle\nabla G_{\beta}^{\delta}\ast\theta,(K_{\beta}\ast\theta^{1})\theta\rangle\,|\\ &\leq\|\theta\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|(K_{\beta}\ast\theta^{1})\nabla G_{\beta}^{\delta}\ast\theta\|_{\dot{H}^{\frac{\beta}{2}-1+\alpha+\epsilon}}\\ &\lesssim\|\theta\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|K_{\beta}\ast\theta^{1}\|_{\dot{H}^{2(\alpha+\epsilon)}}\|{\nabla G_{\beta}^{\delta}\ast\theta}\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}\\ &\lesssim\|\theta\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}^{2}\|\theta^{1}\|_{\dot{H}^{2(\alpha+\epsilon)-\beta+1}}.\end{split}$
@@
$\langle\nabla G_{\beta}\ast\theta,(K_{\beta}\ast\theta)\theta^{2}\rangle=\langle\nabla(G_{\beta}\ast\theta),\nabla^{\perp}(G_{\beta}\ast\theta)\,\theta^{2}\rangle=0,$
@@
$\begin{split}|I_{2}|&=|\langle\nabla(G_{\beta}^{\delta}-G_{\beta})\ast\theta,(K_{\beta}*\theta)\theta^{2}\rangle|\\ &\leq\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|(K_{\beta}*\theta)\nabla(G_{\beta}^{\delta}-G_{\beta})\ast\theta\|_{\dot{H}^{\frac{\beta}{2}-1+\alpha+\epsilon}}\\ &\lesssim\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|K_{\beta}\ast\theta\|_{\dot{H}^{2(\alpha+\epsilon)}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})\ast\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}\\ &\leq\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|\theta\|_{\dot{H}^{2(\alpha+\epsilon)-\beta+1}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})\ast\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}.\end{split}$
@@
$\begin{split}|I_{2}|&\lesssim\big{(}\|\theta^{1}\|_{L^{\tilde{p}}}+\|\theta^{2}\|_{L^{\tilde{p}}}\big{)}\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})\ast\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}\\ &\lesssim(C_{1}+C_{2})\|\theta^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}\|\nabla(G_{\beta}^{\delta}-G_{\beta})\ast\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}.\end{split}$
@@
$\int_{0}^{T}\mathbb{E}[|I_{2}|]\,\mathrm{d}r\lesssim\bigg{(}\int_{0}^{T}\mathbb{E}\Big{[}\|\theta_{r}^{2}\|_{\dot{H}^{-\frac{\beta}{2}+1-\alpha-\epsilon}}^{2}\Big{]}\mathrm{d}r\bigg{)}^{1/2}\bigg{(}\int_{0}^{T}\mathbb{E}\Big{[}\|\nabla(G_{\beta}^{\delta}-G_{\beta})\ast\theta\|_{\dot{H}^{\frac{\beta}{2}-\alpha-\epsilon}}^{2}\Big{]}\mathrm{d}r\bigg{)}^{1/2}.$
@@
For $(t,x)\in(0,\infty)\times\mathbb{R}^{d}$, define $q(t,x)=t(t^{2/\beta}+|x|^{2})^{-\frac{d+\beta}{2}}$. The following bound of the kernel $p(t,x)$ is well known, see for example [6; 32].
@@
From a probabilistic point of view, the fractional Laplacian $-(-\Delta)^{\frac{\beta}{2}}$ is the infinitesimal generator of the $\beta$-stable process, which is a pure jump Markov process. That is to say, $p(t,x)$ is the transition probability of a $\beta$-stable process $Z=(Z_{t})_{t\geq 0}$. Then by the Bochner subordination formula (see e.g. [5, Proposition 8.6]), we have
@@
where $h(s,x)=(4\pi s)^{-\frac{d}{2}}e^{-\frac{|x|^{2}}{4s}}$ is the heat kernel in $\mathbb{R}^{d}$, $\rho_{t}$ is the probability density function of $S_{t}$. Accroding to [6, Lemma 5], there exists a constant $c$ such that $\rho_{1}(s)\leq c\,s^{-1-\frac{\beta}{2}}$ for all $s>0$, so by scaling property $\rho_{t}(s)=t^{-\frac{2}{\beta}}\rho_{1}(t^{-\frac{2}{\beta}}s)$, we have
@@ Appendix B Estimating the remainders
The proofs are almost identical to [11, Lemma A.1], so we omit the details here.
@@
Since $21<p<\frac{2}{3-\beta+\gamma}<$ 2, we can choose $0<\lambda<1-\gamma$ such that $22+(\gamma+\lambda-1)p<$ 2 and $p(4-\beta-\lambda)<2$, making the right-hand side above integrable. Therefore, $g$ has finite $W^{1+\gamma,p}$ norm for $10<\gamma<\beta-$ 1 and $21<p<\frac{2}{3-\beta+\gamma}<$ 2. Then we have (see [41, Subsection 2.5])